Interactive LearningWare 30.1 reviews the concepts that play roles in this problem. A hydrogen atom emits a photon that has momentum 6.460 × 10-27 kg·m/s. This photon is emitted because the electron in the atom falls from a higher energy level into the n = 1 level. What is the quantum number of the level from which the electron falls? Use values of h = 6.626 × 10-34 J·s, c = 2.998 × 108 m/s, and e = 1.602 × 10-19 C.
Answer #1

Momentum, p = 6.46*10-27 kg.m/s

energy, E = pc = 6.46*10-27 * 3*108 = 1.938*10-18 J = 12.1125 eV

E = -13.6(1/n2 - 1/12)

1 - 1/n2 = 0.8906

n = 3

electron falls from n = 3 level.